# Bayesian Optimization for Non-Convex Two-Stage Stochastic Optimization Problems

### v1

@@ A.4 Convergence to optimal recommendations
$\mathrm{Var}\Bigl{[}\bigl{(}\overline{f}({\bm{x}},{\bm{g}})-\overline{\mu}^{\infty}({\bm{x}},{\bm{g}})\bigr{)}-\bigl{(}\overline{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\overline{\mu}^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime})\bigr{)}\,\Big{|}\,\mathcal{F}^{\infty}\Bigr{]}\\ =\underbrace{k^{\infty}({\bm{x}},{\bm{g}};\,{\bm{x}},{\bm{g}})-k^{\infty}({\bm{x}},{\bm{g}};\,{\bm{x}}^{\prime},{\bm{g}}^{\prime})}_{0\text{ a.s.}}+\underbrace{k^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime};\,{\bm{x}}^{\prime},{\bm{g}}^{\prime})-k^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime};\,{\bm{x}},{\bm{g}})}_{0\text{ a.s.}}=0$
almost surely by Proposition A.14. Also, $\mathbb{E}[(\overline{f}({\bm{x}},{\bm{g}})-\overline{\mu}^{\infty}({\bm{x}},{\bm{g}}))-(\overline{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\overline{\mu}^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime}))\,|\,\mathcal{F}^{\infty}]=0$ almost surely. Write ${\textnormal{z}}=(\overline{f}({\bm{x}},{\bm{g}})-\overline{\mu}^{\infty}({\bm{x}},{\bm{g}}))-(\overline{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\overline{\mu}^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime}))$. The law of total variance gives
@@
Therefore, z is a random variable with zero mean and zero variance, and so is almost surely zero. That is, $\overline{f}({\bm{x}},{\bm{g}})-\overline{\mu}^{\infty}({\bm{x}},{\bm{g}})=\overline{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\overline{\mu}^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime})$ almost surely. Taking ${\textnormal{c}}_{2}=\overline{f}({\bm{x}},{\bm{g}})-\overline{\mu}^{\infty}({\bm{x}},{\bm{g}})$ completes the proof. ∎
Having established that $\overline{f}$ and $\overline{\mu}^{\infty}$ are pointwise almost surely equal up to a constant, we now extend this to show that this holds for all points simultaneously. That is, that their sample paths are almost surely equal up to a constant.
@@
$\mathbb{P}\biggl{(}\forall\,{\bm{x}}\in{\mathbb{X}}\;\forall\,{\bm{g}}:{\mathbb{U}}\to{\mathbb{Y}}\text{ measurable}\qquad\overline{f}({\bm{x}},{\bm{g}})=\overline{\mu}^{\infty}({\bm{x}},{\bm{g}})+{\textnormal{c}}_{2}\biggr{)}=1.$
@@
$\mathbb{P}\biggl{(}\forall\,{\bm{x}}\in{\mathbb{X}}\cap\mathbb{Q}^{d_{x}}\;\forall\,{\bm{g}}:{\mathbb{U}}\to{\mathbb{Y}}\cap\mathbb{Q}^{d_{y}}\text{ simple}\qquad\overline{f}({\bm{x}},{\bm{g}})=\overline{\mu}^{\infty}({\bm{x}},{\bm{g}})+{\textnormal{c}}_{2}\biggr{)}=1.$ (31)
@@
$\underbrace{\mathbb{E}_{\mathbf{u}}\Bigl{[}f({\bm{x}}^{i},{\bm{g}}^{i}({\bm{u}}),{\bm{u}})-\mu^{\infty}({\bm{x}}^{i},{\bm{g}}^{i}({\bm{u}}),{\bm{u}})\Bigr{]}}_{\overline{f}({\bm{x}}^{i},{\bm{g}}^{i})-\overline{\mu}^{\infty}({\bm{x}},{\bm{g}}^{i})}\to\underbrace{\mathbb{E}_{\mathbf{u}}\Bigl{[}f({\bm{x}},{\bm{g}}({\bm{u}}),{\bm{u}})-\mu^{\infty}({\bm{x}},{\bm{g}}({\bm{u}}),{\bm{u}})\Bigr{]}}_{\overline{f}({\bm{x}},{\bm{g}})-\overline{\mu}^{\infty}({\bm{x}},{\bm{g}})}.$
On the event $A$ considered in Equation 31, each term in the sequence on the left hand side is equal to ${\textnormal{c}}_{2}$, and therefore the limit on the right must equal ${\textnormal{c}}_{2}$ as well. Whence, on $A$, $\overline{f}({\bm{x}},{\bm{g}})-\overline{\mu}^{\infty}({\bm{x}},{\bm{g}})={\textnormal{c}}_{2}$. Our choice of ${\bm{x}}$ and ${\bm{g}}$ was arbitrary, and so
$\mathbb{P}\biggl{(}\forall\,{\bm{x}}\in{\mathbb{X}}\;\forall\,{\bm{g}}:{\mathbb{U}}\to{\mathbb{Y}}\text{ measurable}\qquad\overline{f}({\bm{x}},{\bm{g}})=\overline{\mu}^{\infty}({\bm{x}},{\bm{g}})+{\textnormal{c}}_{2}\biggr{)}=1.$
@@
$\overline{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\to\sup_{\begin{subarray}{c}{\mathbf{x}}\in{\mathbb{X}}\\ {\mathbf{g}}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\overline{f}({\mathbf{x}},{\mathbf{g}})\qquad\text{as}\qquad n\to\infty$
@@
By Lemma A.2, $\overline{f}$ is bounded and attains its bounds, so let ${\mathbf{x}}^{*},{\mathbf{g}}^{*}\in\operatorname*{arg\,max}_{{\mathbf{x}},{\mathbf{g}}}\overline{f}({\mathbf{x}},{\mathbf{g}})$ be random elements maximizing $\overline{f}$. For any $n$, $\overline{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\leq\sup_{{\mathbf{x}},{\mathbf{g}}}\overline{f}({\mathbf{x}},{\mathbf{g}})$ almost surely. Therefore, it suffices to prove that $\limsup_{n\to\infty}\overline{f}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\overline{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})=0$.
@@
$\limsup_{n\to\infty}$ $\;\overline{f}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\overline{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})$
$\leq$ $\limsup_{n\to\infty}\Bigl{(}\overline{f}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\overline{\mu}^{\infty}({\mathbf{x}}^{*},{\mathbf{g}}^{*})\Bigr{)}-\Bigl{(}\overline{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})-\overline{\mu}^{\infty}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\Bigr{)}$
$+$ $\limsup_{n\to\infty}\Bigl{(}\overline{\mu}^{\infty}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\overline{\mu}^{n}({\mathbf{x}}^{*},{\mathbf{g}}^{*})\Bigr{)}-\Bigl{(}\overline{\mu}^{\infty}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})-\overline{\mu}^{n}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\Bigr{)}$
$+$ $\limsup_{n\to\infty}\overline{\mu}^{n}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\overline{\mu}^{n}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n}).$
@@
$\Bigl{(}\overline{f}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\overline{\mu}^{\infty}({\mathbf{x}}^{*},{\mathbf{g}}^{*})\Bigr{)}-\Bigl{(}\overline{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})-\overline{\mu}^{\infty}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\Bigr{)}={\textnormal{c}}_{2}-{\textnormal{c}}_{2}=0$
@@
Since $\overline{\mu}^{n}\to\overline{\mu}^{\infty}$ uniformly, almost surely, as $n\to\infty$ (Proposition A.12), the second line is also almost surely zero. Finally, the third line line is at most zero by optimality of ${\mathbf{x}}^{*n},{\mathbf{g}}^{*n}$. Therefore,
$\limsup_{n\to\infty}\overline{f}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\overline{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\leq 0$
and since $\overline{f}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\overline{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\geq 0$ almost surely as well, we conclude that $\overline{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\to\overline{f}({\mathbf{x}}^{*},{\mathbf{g}}^{*})$ almost surely as $n\to\infty$.
To establish convergence in mean, we appeal to the Dominated Convergence Theorem with dominating variable $\sup_{{\mathbf{x}},{\mathbf{g}}}|\overline{f}({\mathbf{x}},{\mathbf{g}})|$. Indeed, this has finite expectation by Lemma A.13. ∎
@@ Appendix B Caching of the covariances
$k^{n}({\bm{x}},{\bm{x}})=k({\bm{x}},{\bm{x}})-{\bm{k}}({\bm{x}},{\bm{X}}^{n})\bigl{(}{\bm{K}}({\bm{X}}^{1:n},{\bm{X}}^{1:n})+\sigma^{2}{\bm{I}}\bigr{)}^{-1}{\bm{k}}({\bm{X}}^{1:n},{\bm{x}}),$ (34)
where ${\bm{X}}^{1:n}$ is a matrix formed by stacking ${\bm{x}}^{1},\dots,{\bm{x}}^{n}$ in the rows, $k$ is the prior covariance function, and ${\bm{k}}$ and ${\bm{K}}$ are vector and matrix valued functions obtained from $k$ by row-wise evaluation. Note that ${\bm{k}}({\bm{x}},{\bm{X}}^{1:n})$ is a row vector while ${\bm{k}}({\bm{X}}^{1:n},{\bm{x}})$ is a column vector. The inversion $({\bm{K}}({\bm{X}}^{1:n},{\bm{X}}^{1:n})+\sigma^{2}{\bm{I}})^{-1}$ is potentially costly, and does not depend on the test input ${\bm{x}}$. It is cached by default in GPyTorch.
@@
$\mu^{n}({\bm{x}}^{\prime})$ $=\mu({\bm{x}}^{\prime})+{\bm{k}}({\bm{x}}^{\prime},{\bm{X}}^{1:n})\bigl{(}{\bm{K}}({\bm{X}}^{1:n},{\bm{X}}^{1:n})+\sigma^{2}{\bm{I}}\bigr{)}^{-1}({\bm{v}}^{1:n}-{\bm{\mu}}({\bm{X}}^{1:n})),$ (35)
$k^{n}({\bm{x}}^{\prime},{\bm{x}})$ $=k({\bm{x}}^{\prime},{\bm{x}})-{\bm{k}}({\bm{x}}^{\prime},{\bm{X}}^{1:n})\bigl{(}{\bm{K}}({\bm{X}}^{1:n},{\bm{X}}^{1:n})+\sigma^{2}{\bm{I}}\bigr{)}^{-1}{\bm{k}}({\bm{X}}^{1:n},{\bm{x}}).$ (36)
Here, $\mu$ is the prior mean, ${\bm{\mu}}$ is the vector-valued function obtained from $\mu$ by row-wise evaluation, and ${\bm{v}}^{1:n}$ is the vector of the first $n$ noisy observations of $f$. As already mentioned, the inversion $({\bm{K}}({\bm{X}}^{1:n},{\bm{X}}^{1:n})+\sigma^{2}{\bm{I}})^{-1}$ does not depend on ${\bm{x}}$ and so should be cached. In fact, for the computation of $\mu^{n}({\bm{x}}^{\prime})$, the whole product $({\bm{K}}({\bm{X}}^{1:n},{\bm{X}}^{1:n})+\sigma^{2}{\bm{I}})^{-1}({\bm{v}}^{1:n}-{\bm{\mu}}({\bm{X}}^{1:n}))$ is cached.
Further, the cross-covariances ${\bm{k}}({\bm{x}}^{\prime},{\bm{X}}^{1:n})$ also do not depend on ${\bm{x}}$ and are computed for many ${\bm{x}}^{\prime}$. Repeating this computation every time the query location ${\bm{x}}$ is updated is costly and wasteful. Writing ${\bm{X}}^{\prime}$ for the matrix of a discretization points ${\bm{x}}^{\prime}\in{{\mathbb{X}}_{\mathrm{disc}}}$, we may cache the matrix ${\bm{K}}({\bm{X}}^{\prime},{\bm{X}}^{1:n})$ during the optimization of $\hat{\alpha}_{\text{KG}}^{n}({\bm{x}})$. Due to the design of the API, this is not done by default in GPyTorch.
In two-stage problems modelled by a GP $f:\tilde{{\mathbb{X}}}\to\mathbb{R}$ with $\tilde{{\mathbb{X}}}={\mathbb{X}}\times{\mathbb{Y}}\times{\mathbb{U}}$, the quantities to be cached are simply
$\bigl{(}{\bm{K}}(\tilde{{\bm{X}}}^{1:n},\tilde{{\bm{X}}}^{1:n})+\sigma^{2}{\bm{I}}\bigr{)}^{-1},$ (37a)
$\bigl{(}{\bm{K}}(\tilde{{\bm{X}}}^{1:n},\tilde{{\bm{X}}}^{1:n})+\sigma^{2}{\bm{I}}\bigr{)}^{-1}({\bm{v}}^{1:n}-{\bm{\mu}}(\tilde{{\bm{X}}}^{1:n})),$ (37b)
${\bm{K}}(\tilde{{\bm{X}}}^{\prime},\tilde{{\bm{X}}}^{1:n}).$ (37c)
In the experiments in this paper, these quantities were cached in the implementation of the joint and alternating knowledge gradient acquisition functions.

### v2

@@ A.4 Convergence to optimal recommendations
$\mathrm{Var}\Bigl{[}\bigl{(}\bar{f}({\bm{x}},{\bm{g}})-\bar{\mu}^{\infty}({\bm{x}},{\bm{g}})\bigr{)}-\bigl{(}\bar{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\bar{\mu}^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime})\bigr{)}\,\Big{|}\,\mathcal{F}^{\infty}\Bigr{]}\\ =\underbrace{k^{\infty}({\bm{x}},{\bm{g}};\,{\bm{x}},{\bm{g}})-k^{\infty}({\bm{x}},{\bm{g}};\,{\bm{x}}^{\prime},{\bm{g}}^{\prime})}_{0\text{ a.s.}}+\underbrace{k^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime};\,{\bm{x}}^{\prime},{\bm{g}}^{\prime})-k^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime};\,{\bm{x}},{\bm{g}})}_{0\text{ a.s.}}=0$
almost surely by Proposition A.14. Also, $\mathbb{E}[(\bar{f}({\bm{x}},{\bm{g}})-\bar{\mu}^{\infty}({\bm{x}},{\bm{g}}))-(\bar{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\bar{\mu}^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime}))\,|\,\mathcal{F}^{\infty}]=0$ almost surely. Write ${\textnormal{z}}=(\bar{f}({\bm{x}},{\bm{g}})-\bar{\mu}^{\infty}({\bm{x}},{\bm{g}}))-(\bar{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\bar{\mu}^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime}))$. The law of total variance gives
@@
Therefore, z is a random variable with zero mean and zero variance, and so is almost surely zero. That is, $\bar{f}({\bm{x}},{\bm{g}})-\bar{\mu}^{\infty}({\bm{x}},{\bm{g}})=\bar{f}({\bm{x}}^{\prime},{\bm{g}}^{\prime})-\bar{\mu}^{\infty}({\bm{x}}^{\prime},{\bm{g}}^{\prime})$ almost surely. Taking ${\textnormal{c}}_{2}=\bar{f}({\bm{x}},{\bm{g}})-\bar{\mu}^{\infty}({\bm{x}},{\bm{g}})$ completes the proof. ∎
Having established that $\bar{f}$ and $\bar{\mu}^{\infty}$ are pointwise almost surely equal up to a constant, we now extend this to show that this holds for all points simultaneously. That is, that their sample paths are almost surely equal up to a constant.
@@
$\mathbb{P}\biggl{(}\forall\,{\bm{x}}\in{\mathbb{X}}\;\forall\,{\bm{g}}:{\mathbb{U}}\to{\mathbb{Y}}\text{ measurable}\qquad\bar{f}({\bm{x}},{\bm{g}})=\bar{\mu}^{\infty}({\bm{x}},{\bm{g}})+{\textnormal{c}}_{2}\biggr{)}=1.$
@@
$\mathbb{P}\biggl{(}\forall\,{\bm{x}}\in{\mathbb{X}}\cap\mathbb{Q}^{d_{x}}\;\forall\,{\bm{g}}:{\mathbb{U}}\to{\mathbb{Y}}\cap\mathbb{Q}^{d_{y}}\text{ simple}\qquad\bar{f}({\bm{x}},{\bm{g}})=\bar{\mu}^{\infty}({\bm{x}},{\bm{g}})+{\textnormal{c}}_{2}\biggr{)}=1.$ (31)
@@
$\underbrace{\mathbb{E}_{\mathbf{u}}\Bigl{[}f({\bm{x}}^{i},{\bm{g}}^{i}({\bm{u}}),{\bm{u}})-\mu^{\infty}({\bm{x}}^{i},{\bm{g}}^{i}({\bm{u}}),{\bm{u}})\Bigr{]}}_{\bar{f}({\bm{x}}^{i},{\bm{g}}^{i})-\bar{\mu}^{\infty}({\bm{x}},{\bm{g}}^{i})}\to\underbrace{\mathbb{E}_{\mathbf{u}}\Bigl{[}f({\bm{x}},{\bm{g}}({\bm{u}}),{\bm{u}})-\mu^{\infty}({\bm{x}},{\bm{g}}({\bm{u}}),{\bm{u}})\Bigr{]}}_{\bar{f}({\bm{x}},{\bm{g}})-\bar{\mu}^{\infty}({\bm{x}},{\bm{g}})}.$
On the event $A$ considered in Equation 31, each term in the sequence on the left hand side is equal to ${\textnormal{c}}_{2}$, and therefore the limit on the right must equal ${\textnormal{c}}_{2}$ as well. Whence, on $A$, $\bar{f}({\bm{x}},{\bm{g}})-\bar{\mu}^{\infty}({\bm{x}},{\bm{g}})={\textnormal{c}}_{2}$. Our choice of ${\bm{x}}$ and ${\bm{g}}$ was arbitrary, and so
$\mathbb{P}\biggl{(}\forall\,{\bm{x}}\in{\mathbb{X}}\;\forall\,{\bm{g}}:{\mathbb{U}}\to{\mathbb{Y}}\text{ measurable}\qquad\bar{f}({\bm{x}},{\bm{g}})=\bar{\mu}^{\infty}({\bm{x}},{\bm{g}})+{\textnormal{c}}_{2}\biggr{)}=1.$
@@
$\bar{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\to\sup_{\begin{subarray}{c}{\mathbf{x}}\in{\mathbb{X}}\\ {\mathbf{g}}:{\mathbb{U}}\to{\mathbb{Y}}\end{subarray}}\bar{f}({\mathbf{x}},{\mathbf{g}})\qquad\text{as}\qquad n\to\infty$
@@
By Lemma A.2, $\bar{f}$ is bounded and attains its bounds, so let ${\mathbf{x}}^{*},{\mathbf{g}}^{*}\in\operatorname*{arg\,max}_{{\mathbf{x}},{\mathbf{g}}}\bar{f}({\mathbf{x}},{\mathbf{g}})$ be random elements maximizing $\bar{f}$. For any $n$, $\bar{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\leq\sup_{{\mathbf{x}},{\mathbf{g}}}\bar{f}({\mathbf{x}},{\mathbf{g}})$ almost surely. Therefore, it suffices to prove that $\limsup_{n\to\infty}\bar{f}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\bar{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})=0$.
@@
$\limsup_{n\to\infty}$ $\;\bar{f}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\bar{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})$
$\leq$ $\limsup_{n\to\infty}\Bigl{(}\bar{f}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\bar{\mu}^{\infty}({\mathbf{x}}^{*},{\mathbf{g}}^{*})\Bigr{)}-\Bigl{(}\bar{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})-\bar{\mu}^{\infty}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\Bigr{)}$
$+$ $\limsup_{n\to\infty}\Bigl{(}\bar{\mu}^{\infty}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\bar{\mu}^{n}({\mathbf{x}}^{*},{\mathbf{g}}^{*})\Bigr{)}-\Bigl{(}\bar{\mu}^{\infty}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})-\bar{\mu}^{n}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\Bigr{)}$
$+$ $\limsup_{n\to\infty}\bar{\mu}^{n}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\bar{\mu}^{n}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n}).$
@@
$\Bigl{(}\bar{f}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\bar{\mu}^{\infty}({\mathbf{x}}^{*},{\mathbf{g}}^{*})\Bigr{)}-\Bigl{(}\bar{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})-\bar{\mu}^{\infty}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\Bigr{)}={\textnormal{c}}_{2}-{\textnormal{c}}_{2}=0$
@@
Since $\bar{\mu}^{n}\to\bar{\mu}^{\infty}$ uniformly, almost surely, as $n\to\infty$ (Proposition A.12), the second line is also almost surely zero. Finally, the third line line is at most zero by optimality of ${\mathbf{x}}^{*n},{\mathbf{g}}^{*n}$. Therefore,
$\limsup_{n\to\infty}\bar{f}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\bar{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\leq 0$
and since $\bar{f}({\mathbf{x}}^{*},{\mathbf{g}}^{*})-\bar{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\geq 0$ almost surely as well, we conclude that $\bar{f}({\mathbf{x}}^{*n},{\mathbf{g}}^{*n})\to\bar{f}({\mathbf{x}}^{*},{\mathbf{g}}^{*})$ almost surely as $n\to\infty$.
To establish convergence in mean, we appeal to the Dominated Convergence Theorem with dominating variable $\sup_{{\mathbf{x}},{\mathbf{g}}}|\bar{f}({\mathbf{x}},{\mathbf{g}})|$. Indeed, this has finite expectation by Lemma A.13. ∎
@@ Appendix B Caching of the covariances
$k^{n}({\bm{x}},{\bm{x}})=k({\bm{x}},{\bm{x}})-{\bm{k}}_{\bm{X}}({\bm{x}})^{T}\bigl{(}{\bm{K}}_{{\bm{X}}{\bm{X}}}+\sigma^{2}{\bm{I}}\bigr{)}^{-1}{\bm{k}}_{\bm{X}}({\bm{x}}),$ (34)
where ${\bm{k}}_{\bm{X}}({\bm{x}})$ is the column vector with entries $[{\bm{k}}_{\bm{X}}({\bm{x}})]_{i}=k({\bm{x}}^{i},{\bm{x}})$, and where ${\bm{K}}_{{\bm{X}}{\bm{X}}}$ is the matrix with elements $[{\bm{K}}_{{\bm{X}}{\bm{X}}}]_{ij}=k({\bm{x}}^{i},{\bm{x}}^{j})$. Recall ${\bm{x}}^{1},\dots,{\bm{x}}^{n}\in{\mathbb{X}}$ are the observation locations. The inversion $({\bm{K}}_{{\bm{X}}{\bm{X}}}+\sigma^{2}{\bm{I}})^{-1}$ is potentially costly, and does not depend on the test input ${\bm{x}}$. It is cached by default in GPyTorch.
@@
$\mu^{n}({\bm{x}}^{\prime})$ $=\mu({\bm{x}}^{\prime})+{\bm{k}}_{\bm{X}}({\bm{x}}^{\prime})^{T}\bigl{(}{\bm{K}}_{{\bm{X}}{\bm{X}}}+\sigma^{2}{\bm{I}}\bigr{)}^{-1}({\bm{v}}-{\bm{\mu}}_{\bm{X}}),$ (35)
$k^{n}({\bm{x}}^{\prime},{\bm{x}})$ $=k({\bm{x}}^{\prime},{\bm{x}})-{\bm{k}}_{\bm{X}}({\bm{x}}^{\prime})^{T}\bigl{(}{\bm{K}}_{{\bm{X}}{\bm{X}}}+\sigma^{2}{\bm{I}}\bigr{)}^{-1}{\bm{k}}_{\bm{X}}({\bm{x}}).$ (36)
Here, ${\bm{\mu}}_{\bm{X}}=(\mu({\bm{x}}^{1}),\dots,\mu({\bm{x}}^{n}))$ is the column vector of prior means and ${\bm{v}}=({v}^{1},\dots,{v}^{n})$ is the vector of noisy observations of $f$. As already mentioned, the inversion $({\bm{K}}_{{\bm{X}}{\bm{X}}}+\sigma^{2}{\bm{I}})^{-1}$ does not depend on ${\bm{x}}$ and so should be cached. In fact, for the computation of $\mu^{n}({\bm{x}}^{\prime})$, the whole product $({\bm{K}}_{{\bm{X}}{\bm{X}}}+\sigma^{2}{\bm{I}})^{-1}({\bm{v}}-{\bm{\mu}}_{\bm{X}})$ is cached.
Further, the cross-covariances ${\bm{k}}_{\bm{X}}({\bm{x}}^{\prime})$ also do not depend on ${\bm{x}}$ and are computed for many ${\bm{x}}^{\prime}$. Repeating this computation every time the query location ${\bm{x}}$ is updated is costly and wasteful. Writing ${\bm{X}}^{\prime}$ for the matrix of a discretization points ${\bm{x}}^{\prime}\in{{\mathbb{X}}_{\mathrm{disc}}}$, we may cache the matrix ${\bm{K}}_{{\bm{X}}^{\prime}{\bm{X}}}=[k({\bm{x}}^{\prime i},{\bm{x}}^{j})]_{i=1,\dots,N_{x}}^{j=1,\dots,n}$ during the optimization of $\hat{\alpha}_{\text{KG}}^{n}({\bm{x}})$. Due to the design of the API, this is not done by default in GPyTorch.
In two-stage problems modeled by a GP $f:\tilde{{\mathbb{X}}}\to\mathbb{R}$ with $\tilde{{\mathbb{X}}}={\mathbb{X}}\times{\mathbb{Y}}\times{\mathbb{U}}$, the quantities to be cached are simply $\bigl{(}{\bm{K}}_{\tilde{{\bm{X}}}\tilde{{\bm{X}}}}+\sigma^{2}{\bm{I}}\bigr{)}^{-1}$, $\bigl{(}{\bm{K}}_{\tilde{{\bm{X}}}\tilde{{\bm{X}}}}+\sigma^{2}{\bm{I}}\bigr{)}^{-1}({\bm{v}}-{\bm{\mu}}_{\tilde{{\bm{X}}}})$ and ${\bm{K}}_{\tilde{{\bm{X}}}^{\prime}\tilde{{\bm{X}}}}$. In the experiments in this paper, these quantities were cached in the implementation of the joint and alternating knowledge gradient acquisition functions.